OpenStudy (anonymous):

Find the intersection, if any, of the line x=t, y=1+2t, z=3-t and the line x=-3, y=-6+2u, z=3-6u?

OpenStudy (anonymous):

I know it is not parallel so it would have to be either skew, or intersect at a point...

OpenStudy (anonymous):

What's the question? It's confusing. Due you want point of intersection where all those lines or what? Explain what you want and I can help.

OpenStudy (anonymous):

To find out how I can determine whether the lines are skew or if they intersect at a point.

OpenStudy (experimentx):

suppose they instersect y= 1+2t = -6+2u z = 3-6u = 3-t -------------- we get two linear equations 3-6u = 3-t 1+2t = -6+2u --------------------- solving them we get http://www.wolframalpha.com/input/?i=solve+1%2B2t+%3D+-6%2B2u%2C+3-6u+%3D+3-t t = -21/5, u =-7/10 ---------------- we have for x=t and x=3 or x=-21/5 and x=3 so they are not equal points, hence they do not intersect. Note* I'm not quite sure on this

OpenStudy (anonymous):

I think lines are suppose to be skew... 1. x: t=-3 2. y: 1+2t=-6+2u 2t-2u=-7 3. z: 3-t=3-6u -t+6u=0

OpenStudy (anonymous):

1. + 2. t=-3 -t+6u=0 ---------- 6u=-3 u= -1/2

OpenStudy (experimentx):

what's skew??

OpenStudy (anonymous):

Skew lines are two lines that do not intersect but are not parallel...

OpenStudy (experimentx):

Oh ... neither parallel now intersecting right?? you know how to find parametric equation right?? the above method provides that it has no solution for skewness write the parametric equation